# Exploring a tentative link between FRBs and pulsars with broad energy distributions and applications for nearby FRB survey strategies

### v1

@@ 4.1 Bright bursts of various radio sources
Figure 5: Specific energy of the brightest bursts versus their detection rates for FRB 20201124A, FRB 20220912A, FRB 20200120E and the RRATs J1846--0257 and J1854+0306. The data are fitted with a power law relation (red dashed line) using a Bayesian framework, and the shaded region represents the 68% credible interval for the fit. Posterior distributions and a parameter correlation plot for the power law fit parameters are shown in the inset.
Sources with flat energy distributions ($\alpha\gtrsim-1$) at high energies suggest that their “bright pulses” might contribute significantly to the total radio pulse energy release. Therefore, the detection of the brightest bursts from these sources could provide insight into a potential connection between cosmological high-luminosity radio bursts and nearby low-luminosity ones. Figure 5 presents the specific energy of the brightest bursts and the corresponding detection rates for the studied sources with flat energy distributions, i.e., RRATs J1846--0257 and J1854+0306, FRB 20200120E, FRB 20201124A and FRB 20220912A. Interestingly, these data points align well with a simple power-law fit of $\alpha=-0.26\pm 0.05$. Due to the sample size of only five points and the stochastic nature of detecting the “brightest burst”, we remain cautious about interpreting this correlation as a physical relationship. Nevertheless, this trend suggests that high-luminosity radio bursts, such as FRBs, could be naturally derived from low-luminosity radio bursts, such as those from pulsars. Although an extremely shallow power-law index is implied, the flatter energy distributions at higher energies observed in these sources may provide clues for understanding this tentative connection.
@@ 4.2 Survey strategies for nearby FRBs
As presented in Figure 2 and 3 of Section 3, the fitting of the cumulative energy distribution at high energies using a power-law function yields better results with smaller uncertainties compared to the differential distributions. This difference may arise because the limited burst numbers for the studied sources are insufficient to provide a reliable sample for the differential distributions. Therefore, we used the power-law indices ($\alpha$) derived from the cumulative energy distribution to estimate the detection rate in this section, which is given by:
where $N(\geq E)$ represents the predicted number of bursts with specific energy above $E$.
For a telescope operating at a fluence completeness threshold of $F_{\rm det}$, bursts with energy above the threshold $E_{\rm det}=4\pi D_{\rm det}^{2}F_{\rm det}$ are expected to be detected, where $D_{\rm det}$ is the distance to the target source, and assuming a detected burst rate of $R_{\rm det}$. Applying Equation 2, the burst rate expected for bursts detectable by an arbitrary radio telescope with a fluence threshold $F_{\rm th}$, for a source at a distance of $D_{\rm tar}$, can be written as:
$\centering R_{\rm th}\approx R_{\rm det}\times\left(\frac{F_{\rm th}}{F_{\rm det}}\right)^{\alpha}\left(\frac{D_{\rm tar}}{D_{\rm det}}\right)^{2\alpha}.\@add@centering$ (3)
Kremer et al. (2023) presented prospects for detecting FRBs from GC in nearby galaxies using FAST (Jiang et al., 2020) and MeerKAT (Bailes et al., 2020), suggesting that NGC 4486 (M87) could be the best candidate target. However, the fluence threshold ($F_{\rm th}$) of a telescope may be significantly affected by the continuum radio flux densities ($S_{\rm tar}$) from the target galaxies (Suresh et al., 2021), which was not thoroughly discussed in detection rate estimation by Kremer et al. (2023). Taking into account $S_{\rm tar}$ and the telescope’s corresponding system-equivalent flux density (SEFD), the radiometer equation (Lorimer & Kramer, 2004) for estimating the minimum detectable fluence can be written as:
@@
At a given distance, the number of potential sources is proportional to the field of view (FOV) of the telescope. Combining the FOV with Equation 3, the detection rate of the telescope is given by:
@@
Figure 6: Parkes cryoPAF 63 pointings covering more than 10,000 valid GCs. The two most suggested pointings, TP0 (Dec $\sim$ 13 deg) and TP1 (Dec $\sim$ 8 deg), are marked by black and red stars, respectively, with their corresponding sky coverage shown by circles.
@@
Table 2: Properties of the 12 GC systems covered by Parkes cryoPAF pointing TP1 (RA = 186.6 deg; Dec = 8.1 deg). $d_{\rm L}$ and $D_{\rm TP}$ represent the distance from the GC system centre to the earth and the pointing, respectively. References for the galaxy continuum radio flux densities at 1.4 GHz ($S_{\rm tar,1.4GHz}$) are: (a) White & Becker (1992), (b) Brown et al. (2011), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/).
Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm V\_GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
NGC4472 (M49) 187.445 8.000 17.0 0.84 7000 7000 $\sim$ 0.8 (a)
NGC4365 186.118 7.318 23.3 0.92 3246 3246 $<$ 0.1 (b)
NGC4434 186.903 8.154 22.5 0.30 141 141 $<$ 0.1 (c)
VCC-1254 187.521 8.074 16.0 0.91 26 26 $<$ 0.1 (c)
VCC-1107 187.126 7.325 16.0 0.93 25 25 $<$ 0.1 (c)
NGC4464 187.339 8.157 15.8 0.73 25 25 $<$ 0.1 (c)
NGC4318 185.680 8.198 22.1 0.92 18 18 $<$ 0.1 (c)
VCC-747 186.199 8.992 18.5 0.98 16 16 $<$ 0.1 (c)
VCC-1167 187.301 7.879 16.0 0.73 14 14 $<$ 0.1 (c)
VCC-571 185.671 7.950 23.8 0.93 11 11 $<$ 0.1 (c)
VCC-1049 186.978 8.090 16.0 0.37 8 8 $<$ 0.1 (c)
VCC-992 186.828 8.213 16.0 0.25 4 4 $<$ 0.1 (c)
@@
A cryogenic phased array receiver (cryoPAF) is currently being completed and tested on the Parkes telescope. The new cryoPAF is expected to have a system temperature of $<$ 20 K and a gain of $0.735{\rm K/Jy}$, resulting in a corresponding SEFD of $S_{\rm Parkes}=27.21$ Jy. Its wide FOV spans up to 3 square degrees and consists of 72 adjacent beams, covering a frequency range of 700-1950 MHz  [^4] As cryoPAF is still being tested, the accurate parameters of the full-beam system performance are unavailable for now. We used the values referenced from the Parkes cryoPAF introduction by Alex Dunning: https://research.csiro.au/ratechnologies/wp-content/uploads/sites/295/2022/11/PAFAR2022-Dunning-CryoPAF_for_Parkes.pdf. Based on the cryoPAF’s sky coverage, which is modelled as a circle with a radius of approximately 1 deg in this work, and the catalogue of GC systems (Harris et al., 2013) within a relatively close distance ($<$ 40 Mpc), we conducted a grid analysis of the sky with a 0.1 deg step size to estimate the number of GCs covered by the cryoPAF FOV. Notably, about 20% of GCs lie within a 4-arcminute radius of the centre of M87 (Strader et al., 2011), where the radio background is comparable to $S_{\rm Parkes}$ (Perley & Butler, 2017). Given this, the cryoPAF observation should cover the majority of GCs in M87 with adequate sensitivity, and we adopt a conservative estimate that 50% of the GCs are valid targets for Parkes observation (denoted as $N_{\rm V\_GC}$). As shown in Figure 6, among the more than 6,000,000 analysed pointings, 63 could cover more than 10,000 valid GCs, all with Dec near 13 or 8 deg. Pointings TP0 (RA = 187.2 deg; Dec = 13.2 deg) and TP1 (RA = 186.6 deg; Dec = 8.1 deg) cover the largest number of valid GCs in these two Dec ranges. Based on this, we suggest these two pointings as the best targets for searching for FRBs in nearby GCs. As listed in Tables 1 and 2, the Parkes cryoPAF pointing TP0 could cover 24 GC systems, expected to include 21,607 GCs, of which 15,107 are valid targets. TP1 could cover 12 GC systems, expected to include 10,535 GCs, all of which are valid.
@@
Target Instrument $N_{\rm GC}$ $N_{\rm V\_GC}$ $\alpha$ = --0.2 $\alpha$ = --0.5 $\alpha$ = --1.0 $\alpha$ = --1.5 $\alpha$ = --2.0
M60 FAST 4000 1203 0.16(0.01--0.48) 0.35(0.02--1.06) 1.36(0.07--4.08) 5.21(0.27--15.64) 19.98(1.02--59.93)
M60 MeerKAT 4000 4000 0.37(0.02--1.11) 0.49(0.02--1.46) 0.77(0.04--2.31) 1.21(0.06--3.64) 1.92(0.1--5.75)
TP0 Parkes cryoPAF 21607 15107 1.3(0.07--3.89) 1.53(0.08--4.6) 2.03(0.1 --6.1) 2.72(0.14--8.15) 3.65(0.19--10.96)
TP1 Parkes cryoPAF 10535 10535 0.88(0.04--2.65) 1.01(0.05--3.03) 1.28(0.07--3.84) 1.65(0.08--4.94) 2.15(0.11--6.44)
[…]

### v2

@@ 4.1 Bright bursts of various radio sources
Figure 2: Specific energy of the brightest bursts versus their detection rates for FRB 20201124A, FRB 20220912A, FRB 20200120E and the RRATs J1846-0257 and J1854+0306. The data are fitted with a power law relation (red dashed line) using a Bayesian framework, and the shaded region represents the 68% credible interval for the fit. Posterior distributions and a parameter correlation plot for the power law fit parameters are shown in the inset. Three horizontal contour lines are added at 100 hours, 1000 hours, and 10000 hours observation time, with their intersections with each source’s distribution line indicating the expected rate and energy of the brightest bursts detectable for these observation durations.
Sources with flat energy distributions ($\alpha\gtrsim-1$) at high energies suggest that their “bright pulses” might contribute significantly to the total radio pulse energy release. Therefore, the detection of the brightest bursts from these sources could provide insight into a potential connection between cosmological high-luminosity radio bursts and nearby low-luminosity ones. Figure 2 presents the specific energy of the brightest bursts and the corresponding detection rates [^4] Here, we estimate the rate based on a single detected event, with a $1\sigma$ confidence interval from 0.173 to 3.3 (Gehrels, 1986). for the studied sources with flat energy distributions, i.e., RRATs J1846-0257 and J1854+0306, FRB 20200120E, FRB 20201124A and FRB 20220912A. Interestingly, these data points could align with a power-law fit of $\alpha_{\rm Bri}=-0.26\pm 0.05$. Due to the sample size of only five points and the stochastic nature of detecting the “brightest burst”, which usually depends on the telescope sensitivities and monitoring lengths, we will not suggest interpreting this correlation as a physical relationship at this stage. Three horizontal contours at 100, 1000, and 10000 hours mark the expected rate and energy of the brightest detectable bursts at each observation duration. Even extending to 10,000 h, the expected brightest RRAT energies lie below nearby FRBs, and nearby FRBs lie below bright cosmological FRBs. This trend suggests that a much flatter index than that observed in current sources such as FRB 20200120E is required to naturally derive high-luminosity radio bursts, such as FRBs, from low-luminosity bursts, such as those from pulsars.
It is worth noting that the power-law indices used here were derived from cumulative distributions, in which the data points are not independent. As a result, our line fitting may produce biased results and underestimate the associated uncertainties. This limitation might be further addressed through more robust methods, such as the maximum likelihood approach introduced by James et al. (2019), or through future detections of a sufficient number of bright bursts to allow comparison using differential distributions.
@@ 4.2 Survey strategies for nearby FRBs
For a telescope operating at a fluence completeness threshold of $F_{\rm det}$, bursts with energy above the threshold $E_{\rm det}=4\pi D_{\rm det}^{2}F_{\rm det}$ are expected to be detected, where $D_{\rm det}$ is the distance to the target source, and assuming a reference burst rate of $R_{\rm ref}$. Applying Equation 2, the burst rate expected for bursts detectable by an arbitrary radio telescope with a fluence threshold $F_{\rm th}$, for a source at a distance of $D_{\rm tar}$, can be written as:
$\centering R_{\rm th}\approx R_{\rm ref}\times\left(\frac{F_{\rm th}}{F_{\rm det}}\right)^{\alpha}\left(\frac{D_{\rm tar}}{D_{\rm det}}\right)^{2\alpha}.\@add@centering$ (3)
As discussed by Zhang et al. (2024b), current FRB surveys may suffer from selection effects, potentially missing FRBs with relatively low luminosity and narrow widths from nearby GC systems. For example, the current CHIME system only records baseband data for real-time detections with high signal-to-noise ratios (typically between 10 and 12), and at a time resolution of approximately 1 ms (CHIME/FRB Collaboration et al., 2024). Such a system would only be sensitive to a burst similar to the brightest one from FRB 20200120E (with an effective width of $\leq 0.161$ ms and specific luminosity of $\sim 3.0\times 10^{30}$ erg s-1 Hz-1) at distances around $\sim$10 Mpc, and the event rate is extremely low (Zhang et al., 2024b). For more frequent, lower-luminosity bursts from FRB 20200120E, the system is realistically sensitive only to its current distance of $\sim 3$ Mpc.
Kremer et al. (2023) presented prospects for detecting FRBs from GCs in nearby galaxies using FAST and MeerKAT, suggesting that NGC 4486 (M87) could be the best candidate target. However, the fluence threshold ($F_{\rm th}$) of a telescope may be significantly affected by the continuum radio flux densities ($S_{\rm tar}$) from the target galaxies (Suresh et al., 2021), which was not thoroughly discussed in detection rate estimation by Kremer et al. (2023). Taking into account $S_{\rm tar}$ and the telescope’s corresponding system-equivalent flux density (SEFD), the radiometer equation (Lorimer & Kramer, 2004) for estimating the minimum detectable fluence can be written as:
@@
At a given distance, assuming an isotropic source distribution, the number of potential sources is proportional to the field of view (FOV) of the telescope. Combining the FOV with Equation 3, the detection rate of the telescope is given by:
@@
Figure 3: Parkes cryoPAF 1,832 pointings covering more than 7,000 clean GCs, clustered into six groups. The sky coverage of the most recommended pointing in each group, TP0 (RA = 187.2 deg; Dec = 13.2 deg), TP1 (RA = 186.6 deg; Dec = 8.1 deg), TP2 (RA = 225.9 deg; Dec = 1.5 deg), TP3 (RA = 55.0 deg; Dec = -18.6 deg), TP4 (RA = 54.6 deg; Dec = -35.1 deg) and TP5 (RA = 157.4 deg; Dec = -35.5 deg), is indicated by black circle.
@@
Table 2: Properties of the GC systems covered by Parkes cryoPAF pointings: 12 systems by TP1 (RA = 186.6 deg; Dec = 8.1 deg), 3 systems by TP2 (RA = 225.9 deg; Dec = 1.5 deg), 2 systems by TP3 (RA = 55.0 deg; Dec = -18.6 deg), 21 systems by TP4 (RA = 54.6 deg; Dec = -35.1 deg), and 2 systems by TP5 (RA = 157.4 deg; Dec = -35.5 deg). References for the galaxy continuum radio flux densities at 1.4 GHz ($S_{\rm tar,1.4GHz}$) are: (a) White & Becker (1992), (b) Brown et al. (2011), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/), (d) Condon et al. (1998) and (e) Allison et al. (2014).
Pointing Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm C,GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
TP1 NGC4472 (M49) 187.445 8.000 17.0 0.84 7000 7000 $\sim$ 0.8 (a)
NGC4365 186.118 7.318 23.3 0.92 3246 3246 << 0.1 (b)
NGC4434 186.903 8.154 22.5 0.30 141 141 << 0.1 (c)
VCC-1254 187.521 8.074 16.0 0.91 26 26 << 0.1 (c)
VCC-1107 187.126 7.325 16.0 0.93 25 25 << 0.1 (c)
NGC4464 187.339 8.157 15.8 0.73 25 25 << 0.1 (c)
NGC4318 185.680 8.198 22.1 0.92 18 18 << 0.1 (c)
VCC-747 186.199 8.992 18.5 0.98 16 16 << 0.1 (c)
VCC-1167 187.301 7.879 16.0 0.73 14 14 << 0.1 (c)
VCC-571 185.671 7.950 23.8 0.93 11 11 << 0.1 (c)
VCC-1049 186.978 8.090 16.0 0.37 8 8 << 0.1 (c)
VCC-992 186.828 8.213 16.0 0.25 4 4 << 0.1 (c)
TP2 NGC5846 226.622 1.606 24.9 0.73 4670 4670 << 0.1 (b)
NGC5813 225.297 1.702 32.2 0.64 2900 2900 << 0.1 (b)
NGC5845 226.503 1.634 25.9 0.62 180 180 << 0.1 (c)
TP3 NGC1407 55.049 -18.580 28.8 0.05 7000 7000 << 0.1 (d)
NGC1400 54.879 -18.688 26.4 0.14 2960 2960 << 0.1 (b)
TP4 NGC1399 54.621 -35.451 20.7 0.35 6000 6000 $\sim$ 0.6 (e)
NGC1404 54.716 -35.594 20.4 0.50 770 770 << 0.1 (b)
NGC1427 55.581 -35.393 20.5 0.85 500 500 << 0.1 (b)
NGC1380 54.115 -34.976 18.9 0.42 424 424 << 0.1 (b)
NGC1387 54.238 -35.507 19.8 0.50 390 390 << 0.1 (b)
NGC1374 53.819 -35.226 19.6 0.65 360 360 << 0.1 (b)
NGC1379 54.016 -35.441 20.0 0.59 225 225 << 0.1 (b)
NGC1380B 54.287 -35.195 22.0 0.27 156 156 << 0.1 (b)
NGC1375 53.820 -35.266 20.0 0.66 86 86 << 0.1 (c)
NGC1381 54.132 -35.295 18.8 0.43 71 71 << 0.1 (c)
NGC1380A 54.198 -34.740 16.9 0.49 70 70 << 0.1 (c)
ESO-358-G2 53.879 -34.447 20.8 0.88 60 60 << 0.1 (c)
FCC-182 54.226 -35.375 19.6 0.41 59 59 << 0.1 (c)
NGC1389 54.299 -35.746 20.9 0.69 48 48 << 0.1 (c)
NGC1428 55.596 -35.153 20.7 0.82 42 42 << 0.1 (c)
NGC1427A 55.039 -35.625 16.0 0.63 38 38 << 0.1 (c)
ESO-358-G4 54.538 -34.519 20.7 0.58 30 30 << 0.1 (c)
FCC-136 53.623 -35.546 19.7 0.91 25 25 << 0.1 (c)
FCC-254 55.253 -35.743 20.4 0.83 6 6 << 0.1 (c)
FCC-189 54.284 -34.731 20.0 0.45 1 1 << 0.1 (c)
FCC-146 53.798 -35.323 20.0 0.69 1 1 << 0.1 (c)
TP5 NGC3258 157.223 -35.605 32.1 0.18 6000 6000 << 0.1 (b)
NGC3268 157.503 -35.326 34.8 0.19 4750 4750 << 0.1 (b)
@@
A cryogenic phased array receiver (cryoPAF) is currently being completed and tested on the Parkes telescope. The new cryoPAF is expected to have a system temperature of << 20 K and a gain of $0.735{\rm K/Jy}$, resulting in a corresponding SEFD of $S_{\rm Parkes}=27.21$ Jy. Its wide FOV spans up to 3 square degrees and consists of 72 adjacent beams, covering a frequency range of 700-1950 MHz  [^6] As cryoPAF is still being tested, the accurate parameters of the full-beam system performance are unavailable for now. We used the values referenced from the Parkes cryoPAF introduction by Alex Dunning: https://research.csiro.au/ratechnologies/wp-content/uploads/sites/295/2022/11/PAFAR2022-Dunning-CryoPAF_for_Parkes.pdf. Based on the cryoPAF’s sky coverage, which is modelled as a circle with a radius of approximately 1 deg in this work, and the catalogue of GC systems (Harris et al., 2013) within a relatively close distance (<< 40 Mpc), we conducted a grid analysis of the sky with a 0.1 deg step size to estimate the number of GCs covered by the cryoPAF FOV. Notably, about 20% of GCs lie within a 4-arcminute radius of the centre of M87 (Strader et al., 2011), where the radio background is comparable to $S_{\rm Parkes}$ (Perley & Butler, 2017). Given this, the cryoPAF observation should cover the majority of GCs in M87 with adequate sensitivity, and we adopt a conservative estimate that 50% of the GCs are clean targets for Parkes observation (denoted as $N_{\rm C,GC}$). As shown in Figure 3, among the more than 6,000,000 analysed pointings, 1,832 could cover more than 7,000  [^7] We set this threshold based on the estimated number of $N_{\rm C,GC}$ in GC systems within 40 Mpc. clean GCs, clustered into six groups. Based on both the number of $N_{\rm C,GC}$ within the sky coverage and the distance from the pointing to GC system centres, the most recommended pointings for each group are: TP0 (RA = 187.2 deg; Dec = 13.2 deg), TP1 (RA = 186.6 deg; Dec = 8.1 deg), TP2 (RA = 225.9 deg; Dec = 1.5 deg), TP3 (RA = 55.0 deg; Dec = -18.6 deg), TP4 (RA = 54.6 deg; Dec = -35.1 deg) and TP5 (RA = 157.4 deg; Dec = -35.5 deg). Full details of all 1,832 pointings are provided in the Supplementary Materials. As listed in Tables 1 and 2, Parkes cryoPAF pointing TP0 could cover 24 GC systems, expected to include 21,607 GCs, of which 15,107 are clean targets. TP1 covers 12 GC systems with 10,535 $N_{\rm C,GC}$; TP2 covers 3 GC systems with 7,750 $N_{\rm C,GC}$; TP3 covers 2 GC systems with 9,960 $N_{\rm C,GC}$; TP4 covers 21 GC systems with 9,361 $N_{\rm C,GC}$; and TP5 covers 2 GC systems with 10,750 $N_{\rm C,GC}$.
@@
Target Instrument $N_{\rm GC}$ $N_{\rm C,GC}$ $\alpha$ = -0.2 $\alpha$ = -0.5 $\alpha$ = -1.0 $\alpha$ = -1.5 $\alpha$ = -2.0
M60 FAST 4000 1203 0.16(0.01-0.48) 0.35(0.02-1.06) 1.36(0.07-4.08) 5.21(0.27-15.64) 19.98(1.02-59.93)
M60 MeerKAT 4000 4000 0.37(0.02-1.11) 0.49(0.02-1.46) 0.77(0.04-2.31) 1.21(0.06-3.64) 1.92(0.1-5.75)
TP0 Parkes cryoPAF 21607 15107 1.3(0.07-3.89) 1.53(0.08-4.6) 2.03(0.1 -6.1) 2.72(0.14-8.15) 3.65(0.19-10.96)
TP1 Parkes cryoPAF 10535 10535 0.88(0.04-2.65) 1.01(0.05-3.03) 1.28(0.07-3.84) 1.65(0.08-4.94) 2.15(0.11-6.44)
TP2 Parkes cryoPAF 7750 7750 0.56(0.03-1.68) 0.51(0.03-1.54) 0.45(0.02-1.34) 0.40(0.02-1.19) 0.35(0.02-1.06)
TP3 Parkes cryoPAF 9960 9960 0.71(0.04-2.14) 0.64(0.03-1.92) 0.54(0.03-1.61) 0.45(0.02-1.35) 0.38(0.02-1.13)
TP4 Parkes cryoPAF 9361 9361 0.76(0.04-2.28) 0.82(0.04-2.47) 0.94(0.05-2.82) 1.07(0.05-3.22) 1.23(0.06-3.69)
TP5 Parkes cryoPAF 10750 10750 0.72(0.04-2.16) 0.58(0.03-1.75) 0.41(0.02-1.24) 0.29(0.01-0.87) 0.21(0.01-0.62)
Furthermore, we estimated the predicted FRB detection rates for FAST and MeerKAT targeting M60, as well as for Parkes cryoPAF targeting TP0 and TP1, using power-law indices $\alpha=-0.2,\,-0.5,\,-1.0,\,-1.5$ and 2.0-2.0. Most of the essential values used in this calculation are similar to Kremer et al. (2023). The fluence thresholds of FAST and MeerKAT are 0.015 Jy ms (Niu et al., 2021) and 0.09 Jy ms (Bailes et al., 2020). The reference values for the reference rate ($R_{\rm ref}\sim 0.07$ hr-1), fluence threshold ($F_{\rm det}\sim 5$ Jy ms), and source distance ( $D_{\rm det}\sim 3.63$ Mpc) were derived from CHIME’s observations of FRB 20200120E (CHIME/FRB Collaboration et al., 2021; Bhardwaj et al., 2021; Kirsten et al., 2022). The detection of a single FRB repeater by CHIME in the GCs within the M81 region suggests a specific abundance of $N_{\rm obs}\sim 1/900\approx 1.1\times 10^{-3}$ (Bhardwaj et al., 2021; Harris et al., 2013; Kremer et al., 2023), with a 90% confidence range of $(5.6\times 10^{-3},\,3.3\times 10^{-3})$, incorporating the Poisson probability (Lu et al., 2022). For GC systems with $S_{\rm tar,1.4GHz}<0.1$ Jy, which are negligible compared to the SEFDs of all three telescopes, we ignored their effect on the minimum detectable fluence. For GCs with larger $S_{\rm tar,1.4GHz}$, we used the effective minimum detectable fluence calculated using Equation 4.